SHOW THAT THE PLANE AX+BY+CZ+D=0 TOUCHES THE 3X^2-6Y^2+9Z^2+17=0 FIND THE POINT OF CONTACT
1
Expert's answer
2019-10-03T11:55:07-0400

Given hyperboloid

"3x^2-6y^2+9z^2=-17\\dividing\ by\ 9"

"\frac{x^2}{3} - \frac{2y^2}{3} +\frac{ z^2 }{1} = \frac{- 17}{9}"


Tangential plane at any point "(a,b,c)\ is\ given\ by\\"

"\frac{ax}{3}-\frac{2by}{3}+cz=\frac{-17}{9}"


"\frac{-3ax}{17}+\frac{6by}{17}-\frac{9cz}{17}=1\ \ \ \ .......(1)"


Given plane is

"Ax+By+Cz=-D"

Or,

"\frac{-Ax}{D}-\frac{By}{D}-\frac{Cz}{D}=1\ \ \ \ \ .......(2)"

Comparing coefficient of "(1) \ and\ (2)"

We get

"a=\frac{17A}{3D}"

"b=-\frac{17B}{6D}\\c=\frac{17C}{9D}"

So,

Point of contact = "(a,b,c)=(\frac{17A}{3D},-\frac{17B}{6D},\frac{17C}{9D})"